OpenStudy (anonymous):

WILL GIVE A MEDAL Find the hypotenuse of a right triangle whose legs are sqrt 10 and sqrt 15. A. sqrt 5 B. 5 C. 15

OpenStudy (anonymous):

What I have so far is 10^2 + 15^2 = c^2 100 + 225 = c^2 325 = c^2 5 sqrt 13 = c

OpenStudy (anonymous):

So I don't know which answer choice it would be

OpenStudy (triciaal):